# Measuring the Resolved Star Formation Main Sequence in TNG100: Fitting Technique Matters

### v1

@@ 3 Results
Name Resolution (pc) N $\gamma$ $y_{0}$ $\gamma^{\prime}$ $\Sigma_{*}^{\rm to}$ ${\rm R}^{2}$
fid-OLS $\sim 0.75$ kpc $64$ $0.71\pm 0.01$ $-7.51\pm 0.01$ N/A N/A $0.2168$
fid-OLSdoub $\sim 0.75$ kpc $64$ $0.52\pm 0.01$ $-6.27\pm 0.06$ $0.73\pm 0.01$ $6.64\pm 0.02$ $0.2173$
fiducial $\sim 0.75$ kpc $64$ $0.88\pm 0.01$ $-8.15\pm 0.07$ $0.19\pm 0.02$ $7.86\pm 0.02$ $0.9983$
@@
Note. — Fits are done with the ridge line method except for the fid-OLS and fid-OLSdoub cases.
@@ 3.1 Fitting Method
For comparison with other results in the literature, we have performed two OLS fits to star-forming spaxels in our sample of TNG100 galaxies. The first (which we will refer to as fid-OLS) is performed over the same range of stellar masses as the other cases ($10^{6}<\Sigma_{*}/[M_{\odot}{\rm kpc}^{2}]<10^{9}$) and is fit with a single linear function. The fid-OLS fit is represented by the gray line in Figure 1. We also use OLS to fit a double linear function to the full sample of spaxels (fid-OLSdoub) for better comparison to the fits we perform with the ridge line technique. Results for the fits to the rSFMS for our different maps of TNG100 galaxies are summarized in Table 1.
From Table 1, when the rSFMS is fit with a single line, the fid-OLS fit returns a slope that is significantly shallower than those obtained with the ridge line method. The slope obtained with the ridge line method for the fiducial maps is $\gamma=0.88\pm 0.01$, while the slope of the fid-OLS case is $\gamma=0.71\pm 0.01$. The fid-OLSdoub case is interesting because the slope at the low-mass end ($\gamma=0.52\pm 0.01$) is shallower than the slope at the high-mass end ($\gamma^{\prime}=0.73\pm 0.01$, which is at odds with the other results in Table 1.
The significant differences between the OLS fits and the ridge line fit are unsurprising. The main advantage of the ridge line method is that it is not biased by the presence of spaxels transitioning off the main sequence. This transition is expected due to the natural evolution of galaxies.
@@
The $R^{2}$ statistic shows that the OLS models do a poor job of describing the variation of $\Sigma_{\rm SFR}$ with $\Sigma_{*}$. For the fid-OLS and fid-OLSdoub cases, $\rm R^{2}=0.2168$ and $0.2173$, respectively. However, with $\rm R^{2}=0.9983\approx 1$, the ridge line method results in a model that fits the data well. Although it may be possible to improve $\rm R^{2}$ for the OLS cases by imposing additional criteria to limit the sample to spaxels that are actively star-forming, this would defeat the goal of obtaining an objective and broadly comparable rSFMS slope.
@@
In the following subsections, we fit the rSFMS with a double linear function using the ridge line method.
@@ 3.2 Spaxel Resolution
Figure 2: Comparison of the modes of spaxels in the star-forming portion of the $\Sigma_{*}-\Sigma_{\rm SFR}$ diagram derived from maps with different spatial resolutions (left) and different smoothing scales (right). Points represent the ridges identified in each bin of $\Sigma_{*}$ and lines of the same color represent the fit to those ridges. Error bars are approximated as the KDE bandwidth estimated via Scott’s rule.
Previous results (Hani et al., 2020) have indicated that the spatial resolution of the surface density maps may affect the slope obtained for the rSFMS. In Figure 2 (left), we plot the ridge lines from the $2$D distribution in the $\Sigma_{*}-\Sigma_{\rm SFR}$ diagram of star-forming spaxels taken from maps with the following spaxel resolutions: blue: $0.5{\rm kpc}$ ($\sim 0.34h^{-1}{\rm kpc}$), yellow: $\sim 0.74{\rm kpc}$ ($0.5h^{-1}{\rm kpc}$), green: $1{\rm kpc}$ ($\approx 0.68h^{-1}{\rm kpc}$), pink: $\sim 1.5{\rm kpc}$ ($1h^{-1}{\rm kpc}$). Points show the ‘ridge line’ or modes of the distribution of $\Sigma_{\rm SFR}$ in bins of $\Sigma_{*}$, while the solid lines show the fits to the ridge lines. Altering the resolution of the maps affects the surface mass at which the main sequence turns over, but the slopes at the low-mass end agree within $1\sigma$ at all resolutions (see Table 1).
@@
Figure 3 illustrates the effects of spatial resolution on maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ for a particular TNG100 galaxy, TNG100-99-540858222This identifier indicates that this galaxy is from the 99th snapshot of TNG100, and is indexed as 540858 in the subhalo catalog.. TNG100-99-540858 is a main-sequence, central galaxy with a stellar mass of $6.4\times 10^{9}M_{\odot}$. The central star-forming region seen in Figure 3 is spread over a larger area at lower resolutions. This effect contributes to the trend observed in Figure 2 (left).
From Table 1, the value of $\rm R^{2}$ decreases from the fiducial value ($\rm R^{2}=0.9983$) for the res1000 and res1500 cases, which have $\rm R^{2}=0.9866$ and $0.9552$, respectively. This is likely due to the large variation in the ridge line at high masses, which we attribute to a reduction in the number of high-$\Sigma_{*}$ spaxels in these cases. This reduction is a result of there being fewer spaxels in low-resolution maps and that these spaxels have higher surface areas.
@@ 3.3 Adaptive Smoothing Scale
Equation 1 describes how simulated stellar particles are smoothed onto a grid of spaxels to determine spatially resolved star formation rates. In this equation, particles are smoothed over a radius of $h_{sml}$ which is defined as the distance of a particle to its $N^{\rm th}$ nearest neighbor. In Figure 2 (right), we present the rSFMS ridge line that results from maps generated with $N=8$ (pink), $16$ (green), $24$ (yellow), $32$ (blue), and $64$ (purple). From Table 1, the rSFMS slopes agree within $2\sigma$ for all values of $N$ explored here. However, for $\Sigma_{*}<\Sigma_{*}^{\rm to}$, the modes are offset such that as $N$ increases, the zero point decreases. For $\Sigma_{*}>\Sigma_{*}^{\rm to}$, all cases turn over at $\sim\Sigma_{\rm SFR}\sim 10^{-1.25}M_{\odot}{\rm yr}^{-1}{\rm kpc}^{-2}$, but the turnover occurs at higher $\Sigma_{*}$ due to the offset in normalization.
@@ 4 Summary & Discussion
In this Letter, we compared ridge line fits of the rSFMS obtained with ScaleRPy from maps of TNG100 galaxies generated at different spatial resolutions and with different smoothing scales. We find that the ridge line of the rSFMS is fit well by a double linear function (in log space). At the low-mass end, the slopes we obtain with the ridge line method agree with each other within $2\sigma$ in almost all cases. The only exception is the res500 case, where the low-mass slope agrees with all others only at $3\sigma$ [^3] We note that a resolution of $500$ pc is close to the spatial resolution limit of TNG100, which has a gravitational softening length of $\sim 180$ pc. We find that the stellar mass surface density at which the rSFMS turns over ($\Sigma_{*}^{\rm to}$) becomes larger at larger smoothing scales and at finer spatial resolutions. While the normalization of $\Sigma_{\rm SFR}$ at the low-mass end ($y_{0}$) is identical within $1\sigma$ for different spatial resolutions, $y_{0}$ decreases with increasing values of $N$. The rSFMS plateaus at increasing values of $\Sigma_{\rm SFR}$ as spatial resolution becomes finer, but plateaus at roughly the same value of $\Sigma_{\rm SFR}$ across varying smoothing scales.
The cases we present here demonstrate that there are significant differences in the rSFMS from maps generated at different spatial resolutions and with different smoothing scales, particularly at the high-mass end. However, we demonstrate that there is a remarkable consistency in the slope at the low-mass end ($\gamma$). By fitting a double linear function to the ridge line of the rSFMS, we are able to parameterize the differences at the high-mass end, while obtaining an objective slope for the low-mass end of this distribution. Thus, the slope of the low-mass end of the rSFMS obtained with our method is well-suited for comparing this relationship in spatially resolved data obtained with a variety of methods.
As the rSFMS is often fit with OLS and a single linear function in the literature, we fit ‘spaxels’ from our fiducial maps of TNG100 galaxies with OLS. Whether or not we limit the OLS fit to spaxels within the regime where the rSFMS can be approximated as a single linear function, the slope we obtain is significantly lower than the slopes obtained with the ridge line method. This is because the spaxels are not normally-distributed about the main sequence, and spaxels that are actively quenching will bias the distribution toward lower values of $\Sigma_{\rm SFR}$ at higher masses. To obtain reasonable slopes with the OLS method, the fits are often limited to spaxels that are actively star-forming and are within the regime where the rSFMS appears linear. In contrast, the ridge line method does not require pre-selection of star-forming spaxels and only requires limiting the fit to stellar masses where there is enough data to obtain a good kernel density estimate. Thus, the ridge line method allows for the inclusion of more spaxels than the OLS method and does not assume the turnover mass, resulting in a more objective measurement of the slope. This is especially desirable when comparing the rSFMS slope obtained from different observations and simulations, where spatial resolution and sample selection criteria may differ.
Hani et al. (2020) explored the dependence of the rSFMS slope on spatial resolution for galaxies from the FIRE project. They fit the rSFMS with a single linear function and found that the slope became steeper as spatial resolution decreased. While we find that the rSFMS slope at low masses ($\gamma$) is roughly constant with changing spatial resolution, we find that the slope at high masses ($\gamma^{\prime}$) becomes steeper as spatial resolution decreases. The ‘spaxel’ samples used by Hani et al. (2020) appear to be biased toward stellar masses greater than the turnover masses we have identified ($\Sigma_{*}^{\rm to}$), which means the slopes they obtained would be biased toward the high-mass end of the relation. Thus, we find that our results are consistent with those of Hani et al. (2020). This further demonstrates the benefits of fitting the rSFMS with a double linear function. Pessa et al. (2021) fit the rSFMS, resolved molecular gas main sequence, and Kennicutt-Schmidt relation with a single line and found that the slopes were independent of spatial resolution in a spaxel sample drawn from the Physics at High Angular resolutions in Nearby GalaxieS (PHANGS) survey (Leroy et al., 2021). For all three relations, the PHANGS data appear to deviate from linearity at high surface mass densities, although this seems to introduce only a small bias in the slopes obtained with their methodology.
@@
In our efforts to obtain comparable, objective measurements of the rSFMS slope, we have tried to eliminate subjective choices such as imposing bounds on the fit. However, we found it necessary to impose an upper limit on $\Sigma_{*}$, where the identified ridge points are no longer well described by a double linear function. This may be due in part to the relative sparseness of high surface mass density spaxels, resulting from the poor sampling of very high-mass galaxies in the limited volume of TNG100. The advantage of fitting a double linear function is that the slope at the low-mass end ($\gamma$) is not biased by our imposed upper limit. However, this is not the case for the slope at the high-mass end ($\gamma^{\prime}$). It is not clear whether the high-mass end of the rSFMS is best described by a linear function, or what causes the change in slope at $\Sigma_{*}^{\rm to}$. Our results indicate that the location of the turnover point depends at least in part on spatial resolution and the sizes of star-forming regions. Further research will be necessary to understand and constrain the shape of the high-mass end of this distribution and the location of the turnover point.
@@
In conclusion, we find that the slope at the low-mass end of the rSFMS is independent of spatial resolution and smoothing scale when it is measured using a double linear fit to the ridge line of the data. Additionally, this method does not require pre-selection of star-forming spaxels and parameterizes the location at which the rSFMS slope turns over. Thus, this is a promising method for making measurements of scaling relationships that can be compared across studies. The open-source ScaleRPy package facilitates the adoption of the method presented here for both global and resolved scaling relations.

### v3

@@ 3 Results
Name Resolution (pc) N $\gamma$ $y_{0}$ $\gamma^{\prime}$ $\log_{10}[\Sigma_{*}^{\rm to}M_{\odot}^{-1}{\rm kpc}^{2}]$ ${\rm R}^{2}$
fidOLS $\sim 0.75$ kpc $64$ $0.709\pm 0.001$ $-7.509\pm 0.008$ N/A N/A $0.2170$
fidLS-1PL $\sim 0.75$ kpc $64$ $0.554\pm 0.001$ $-5.722\pm 0.009$ N/A N/A $0.1776$
fidLS-2PL $\sim 0.75$ kpc $64$ $0.865\pm 0.006$ $-8.06\pm 0.05$ $0.454\pm 0.002$ $7.738\pm 0.007$ $0.1813$
fidLS2-log $\sim 0.75$ kpc $64$ $0.52\pm 0.01$ $-6.31\pm 0.06$ $0.7323\pm 0.002$ $6.64\pm 0.02$ $0.217$
fiducial $\sim 0.75$ kpc $64$ $0.88\pm 0.01$ $-8.15\pm 0.07$ $0.19\pm 0.02$ $7.86\pm 0.02$ $0.9984$
@@
Note. — Fits are done with the ridge line method except for the fidOLS and fidLS cases.
@@ 3.1 Fitting Method
In this section, we compare results from our ridge line fits to variations of least squares fitting techniques. In Table 1 we report all parameters for the rSFMS fits. We summarize the interesting results for the rSFMS and rKS below, but for precise details of each fit we refer readers to the computational notebook archived on Zenodo22210.5281/zenodo.15047580 (catalog doi: 10.5281/zenodo.15047580); The notebook also includes least squares fits of the rSFMS at other resolutions.
In Figure 1, we compare fits to the rSFMS with the ridge line method to fits obtained with least squares. When the rSFMS is fit with a single power law, the slopes ($0.709\pm 0.001$ for fidOLS and $0.554\pm 0.001$ for LS-1PL) obtained are significantly shallower than that obtained with the ridge line method ($0.88\pm 0.01$) at $\log_{10}\Sigma_{*}<7.86\pm 0.02$ Where least squares are used to fit the rSFMS in $\log_{10}$ space (fidOLS, LS-2PL-log), they are biased toward lower values of $\Sigma_{\rm SFR}$. The fidLS-2PL fit comes closest to identifying the ridge line, with a low-mass slope and intercept in agreement with the ridge line fit. However, the turnover density from fidLS-2PL ($7.738\pm 0.007$) is lower than that from the ridge line ($7.87\pm 0.02$) and the slopes at high-densities also disagree. Although it more closely follows the ridge points (i.e., conditional modes) compared to other least squares fits, the fidLS-2PL method does not result in as good a fit to the ridge points as the ridge line method ($R^{2}=0.6828$ compared to $R^{2}=0.9984$).
@@
The significant differences between the least squares fits and the ridge line fit are unsurprising. For relationships with $\Sigma_{\rm SFR}$, the main advantage of the ridge line method is that it is not biased by the presence of spaxels naturally transitioning off the main sequence. The $R^{2}$ statistic shows that the least squares models do a poor job of describing the variation of $\Sigma_{\rm SFR}$ with $\Sigma_{*}$. For the least squares fits, $R^{2}\sim 0.2$. However, with $\rm R^{2}=0.9984\approx 1$, the ridge line method results in a model that fits the data well. Although it may be possible to improve $\rm R^{2}$ for the least squares cases by imposing additional criteria such as priors and limiting the sample to spaxels that are actively star-forming, this would defeat the goal of obtaining an objective and broadly comparable rSFMS slope.
We now turn to analysis of fits to the rKS and rGMS relations from our fiducial maps. We note that $\Sigma_{\rm gas}$ is the total surface density of all gas, rather than a molecular subset (e.g., $\Sigma_{H_{2}}$).
In Figure 1b, we compare fits to the rKS relation obtained with a double linear ridge line fit and least squares fits. At $\Sigma_{\rm gas}\lesssim 10^{7}M_{\odot}{\rm kpc}^{-2}$, there appears to be essentially no significant relationship between $\Sigma_{\rm gas}$ and $\Sigma_{\rm SFR}$ (the slope of the ridge line fit at $\Sigma_{\rm gas}<10^{7.14\pm 0.03}M_{\odot}{\rm kpc}^{-2}$ is $-0.10\pm 0.08$ and data points are approximately normally distributed over a large range). This is generally consistent with theoretical predictions of a density threshold around $\Sigma_{\rm gas}\sim 10^{7}M_{\odot}{\rm kpc}^{-2}$ that must be met for star formation to occur (e.g., Kennicutt, 1989; Springel & Hernquist, 2003). At higher gas surface densities, the ridge line is fit by a line with a slope of $1.54\pm 0.02$. The slope obtained with least squares range from $1.226\pm 0.002$ (for OLS) to $1.516\pm 0.004$ (for LS-2PL at high $\Sigma_{\rm gas}$). Our analysis of this relationship in TNG100 implies that fits of the rKS to a single power law will be affected by the availability of data for gas at low surface densities. The $R^{2}$ statistic for the ridge line fit to the rKS is $0.9967$, compared to $0.2129$ for the OLS fit.
Unlike the other relations we identify here, the prescriptions to estimate subgrid star formation in the TNG simulations were specifically tuned to reproduce the rKS relation with a slope of $1.5$ at $\Sigma_{\rm gas}\gtrsim 10^{7}\;M_{\odot}\;{\rm kpc}^{2}$ (Springel & Hernquist, 2003). This tuning was not trivial, and the relationship that results from solving the equations which govern star formation in the Springel & Hernquist (2003) model has a slope that is not reported but appears slightly steeper than $1.5$. Because this model is used to calculate the SFR for star-forming gas particles in the TNG100 simulation, any rKS relationship recovered from that simulation should have a slope in agreement with that model. The rKS slope we identify with the ridge line method ($1.54\pm 0.02$), as well as those obtained with least squares fits to double power laws ($1.476\pm 0.002$ and $1.516\pm 0.004$ for LS-2PL and LS-2PL-log, respectively) are generally consistent with a slope of $1.5\pm 0.1$, but we leave a more rigorous statistical comparison for a later work. According to the Springel & Hernquist (2003), the rKS should turn over at $\Sigma_{\rm gas}\sim 10^{7}M_{\odot}{\rm kpc}^{2}M_{\odot}{\rm kpc}^{-2}$. The ridge line method returns a turnover density of $10^{7.14\pm 0.03}$ while LS-2PL-log and LS-2PL return turnover densities of $10^{7.126\pm 0.002}M_{\odot}{\rm kpc}^{-2}$ and $10^{8.389\pm 0.008}M_{\odot}{\rm kpc}^{-2}$, respectively. Thus, only the ridge line method and a least squares fit to a double power law in $\log_{10}$ space returns parameters in agreement with the star formation models built into the TNG100 simulation. The only significant difference between the ridge line fit and LS-2PL-log for the rKS is in the normalization of the relation.
In Figure 1c, we show the distribution of $\Sigma_{*}$ and $\Sigma_{\rm gas}$. Within the bounds of $\Sigma_{*}$ and $\Sigma_{\rm gas}$ that we explore here, fitting the ridge line with a double power law has little effect on the resulting rGMS slope compared to fitting to a single power law, with a difference of $\sim 0.1$ between the slopes at low- and high- density for all double power law fits. The slopes we find from all methods range from $0.27$ to $0.37$. Normalization is the main difference in the rGMS fits done in linear and logarithmic space, as expected. The $R^{2}$ values for the least squares fits to the rGMS are better than those for the least squares fits to the rKS or rSFMS but still objectively poor, with $R^{2}\approx 0.32$ compared to $R^{2}=0.997$ for the ridge line method.
@@
In the following subsections, we fit the rSFMS with a double power law using the ridge line method.
@@ 3.2 Spaxel Resolution
Figure 2: Comparison of the modes of spaxels in the star-forming portion of the $\Sigma_{*}-\Sigma_{\rm SFR}$ diagram derived from maps with different spatial resolutions (left) and different smoothing scales (right). Points represent the ridges identified in each bin of $\Sigma_{*}$ and lines of the same color represent the fit to those ridges. The uncertainties are approximated as the KDE bandwidth estimated via Scott’s rule.
Previous results (Hani et al., 2020) have indicated that the spatial resolution of the surface density maps may affect the slope obtained for the rSFMS. In Figure 2 (left), we plot the ridge lines from the $2$D distribution in the $\Sigma_{*}-\Sigma_{\rm SFR}$ diagram of star-forming spaxels taken from maps with the following spaxel resolutions -- blue: $0.5{\rm kpc}$ ($\sim 0.34h^{-1}{\rm kpc}$), yellow: $\sim 0.74{\rm kpc}$ ($0.5h^{-1}{\rm kpc}$), green: $1{\rm kpc}$ ($\approx 0.68h^{-1}{\rm kpc}$), and pink: $\sim 1.5{\rm kpc}$ ($1h^{-1}{\rm kpc}$). Points show the ‘ridge line’ or modes of the distribution of $\Sigma_{\rm SFR}$ in bins of $\Sigma_{*}$, while the solid lines show the fits to the ridge lines. Altering the resolution of the maps affects the surface mass at which the main sequence turns over, but the slopes at the low-mass end agree within $1\sigma$ at all resolutions (see Table 1).
@@
Figure 3 illustrates the effects of spatial resolution on maps of $\Sigma_{*}$ and $\Sigma_{\rm SFR}$ for a particular TNG100 galaxy, TNG100-99-540858333This identifier indicates that this galaxy is from the 99th snapshot of TNG100, and is indexed as 540858 in the subhalo catalog.. TNG100-99-540858 is a main-sequence, central galaxy with a stellar mass of $6.4\times 10^{9}M_{\odot}$. The central star-forming region seen in Figure 3 is spread over a larger area at lower resolutions. This effect contributes to the trend observed in Figure 2 (left).
From Table 1, the value of $\rm R^{2}$ decreases from the fiducial value ($\rm R^{2}=0.9983$) for the res1000 and res1500 cases, which have $\rm R^{2}=0.9866$ and $0.9552$, respectively. This is likely due to the large variation in the ridge line at high masses, which we attribute to a reduction in the number of high-$\Sigma_{*}$ spaxels in these cases. This reduction is a result of there being fewer spaxels in low-resolution maps and that these spaxels have larger surface areas.
@@ 3.3 Adaptive Smoothing Scale
Equation 1 describes how simulated stellar particles are smoothed onto a grid of spaxels to determine spatially resolved parameters. In this equation, particles are smoothed over a radius of $h_{sml}$ which is defined as the distance of a particle to its $N^{\rm th}$ nearest neighbor.
In Figure 2 (right), we present the rSFMS ridge line that results from maps generated with $N=8$ (pink), $16$ (green), $24$ (yellow), $32$ (blue), and $64$ (purple). From Table 1, the rSFMS slopes agree within $2\sigma$ for all values of $N$ explored here. However, for $\Sigma_{*}<\Sigma_{*}^{\rm to}$, the modes are offset such that as $N$ increases, the zero point decreases. For $\Sigma_{*}>\Sigma_{*}^{\rm to}$, all cases turn over at $\Sigma_{\rm SFR}\sim 10^{-1.25}M_{\odot}{\rm yr}^{-1}{\rm kpc}^{-2}$, but the turnover occurs at higher $\Sigma_{*}$ due to the offset in normalization.
@@ 4 Summary & Discussion
In this work, we have explored different methodologies for obtaining fits to spatially resolved scaling relationships in the TNG100 simulation. We find that fits to the rSFMS and rKS can be highly sensitive to the assumed functional form of the relation and the fitting method. We propose our fiducial method, an extension of the ridge line technique presented by Renzini & Peng (2015), as an objective, consistent approach to measuring scaling relations. We show that applying our ridge line method to the rSFMS for data realized at different spatial resolutions returns low-density slopes that are identical within $1\sigma$. Such insensitivity to spatial resolution is an important criteria for a fitting technqiue that is comparable across many data sets.
An exhaustive exploration of possible functional forms for the rSFMS, rKS, and rGMS is outside the scope of this work. However, we have shown that the rKS and rSFMS are better described by a double, rather than single, power law. Physically, this means that star formation proceeds differently in different density regimes. In the case of the rKS, a transition occurs above some critical density for gravitational stability ($\Sigma_{\rm gas}\gtrsim 10^{7}M_{\odot}{\rm kpc}^{2}$). The physical origin of a bend in the rSFMS is less clear. In Figure 2 we show that the turnover density is dependent on both spatial resolution and smoothing scale. At high-densities, the relationship may not be well described by a power law. However, we demonstrate that there is a remarkable consistency in the slope at the low-density end ($\gamma$). By fitting a double power law to the ridge line of the rSFMS, we are able to parameterize the differences at the high-mass end, while obtaining an objective slope for the low-mass end of this distribution. For the rKS and rSFMS, fits to a single power law consistently underestimate the slope at high- and low- densities, respectively. Better results can be obtained when the least squares fits are limited to a regime that is well described by a single function. However, this introduces a subjective choice into the identification of a scaling relationship slope that limits comparability. Further, selecting a model which parameterizes the location of a regime change, such as the double power law we adopt, allows for more physically significant information to be obtained.
The ridge line method offers significant benefits over straightforward least squares when it comes to identifying a functional form for scaling relations. Identifying the ridge line before fitting a relationship constrains the scope of the problem and minimizes the influence of the secondary processes which drive scatter from the relation. This is evident in the $R^{2}$ statistic for our fits. For the rSFMS, rKS, and rGMS from our fiducial maps $R^{2}>0.99$, while typical $R^{2}$ values for the least squares fits are around $0.2$. Between all of the least squares fits done across the three relationships explored in this work, the best $R^{2}$ value was found for the double linear fit to the rGMS in $\log_{10}$ space ($R^{2}=0.36$). [^4] All $R^{2}$ values and fitting results are available in the computational notebook linked above.
@@
In our efforts to obtain comparable, objective measurements of the rSFMS slope, we have tried to eliminate subjective choices such as imposing bounds on the fit. However, we found it necessary to impose an upper limit on $\Sigma_{*}$, where the identified ridge points are no longer well described by a double power law. This may be due in part to the relative sparseness of high surface mass density spaxels, resulting from the poor sampling of very high-mass galaxies in the limited volume of TNG100. The advantage of fitting a double power law is that the slope at the low-mass end ($\gamma$) is not biased by our imposed upper limit. However, this is not the case for the slope at the high-mass end ($\gamma^{\prime}$). It is not clear whether the high-mass end of the rSFMS is best described by a power law, or what causes the change in slope at $\Sigma_{*}^{\rm to}$. Our results indicate that the location of the turnover point depends at least in part on spatial resolution and the sizes of star-forming regions. Further research will be necessary to understand and constrain the shape of the high-mass end of this distribution and the location of the turnover point.
We identify a systematic offset toward lower normalizations in least squares fits when the relationships are fit in $\log_{10}$ rather than linear space. There are significant benefits to fitting power-law relationships in logarithmic space, given that it linearizes problem and balances the influence of data points that span many orders of magnitude. However, since least squares fits trace the conditional mean of a distribution, fits done in logarithmic space will trace the conditional geometric mean of the distribution. This explains the consistent negative offset we observe when applying least squares to logarithmic data [^5] Mathematically, this is proved by Jensen’s inequality, which states that $\log(E[Y|X])\geq E[\log(Y)|X]$., which was also found by Pessa et al. (2021). Here, we cannot say definitively whether fitting should be performed before or after transformation into logarithmic space. In part, this will depend on whether the scatter from a scaling relationship is normally or log-normally distributed. In some studies, conditional means or medians have been identified in linear space and then fit in logarithmic space (e.g., Medling et al., 2018; Trayford & Schaye, 2019; Pessa et al., 2021), which should mitigate the normalization offset. Modes are invariant to logarithmic transformations and unaffected by asymmetric distributions of data, which may make them a better choice than means or medians. However, we have not directly compared fitting conditional means or medians to the ridge line method, although such a study would be of interest.
Defining scaling relations by the conditional modes is physically motivated. If the goal of obtaining scaling relations or ‘main sequences’ is to describe the typical or ‘main’ way that a galaxy or spatially resolved region behaves, that is described by the mode. In other words, scaling relations should trace the peak of the conditional probability distribution function (i.e., the conditional mode). Our justification for this statement is twofold: [1] scaling relationships arise from fundamental physical processes (e.g., the rGMS arises from gas and stars following the same gravitational potential) while deviations from these relations arises from secondary processes (e.g., star-forming regions moving off the rSFMS due to quenching driven by heating and winds) and [2] scaling relations are often used to estimate an unknown variable from a known variable (e.g., estimating black hole mass from its relationship with central velocity dispersion). Thus, objectively quantifying scaling relationships and their scatter via conditional modes is important for understanding the physical processes that drive them and is necessary for accurately using these relations to estimate unmeasured variables.
As we have shown here, different approaches to fitting scaling relationships make direct comparisons between works challenging, if not impossible. Additionally, for the rKS and rGMS, we have found the relationships only for total gas density, rather than a molecular subset as is more common. However, the slopes we identify with the ridge line technique are in general agreement with the literature. At the low-density end of the rSFMS, we find a slope of $0.88\pm 0.01$ while estimates of the rSFMS slope in the literature range from $\sim 0.5-1$ (e.g., Cano-Díaz et al., 2016; Maragkoudakis et al., 2016; Abdurro’uf & Akiyama, 2017; Liu et al., 2018; Trayford & Schaye, 2019; Erroz-Ferrer et al., 2019; Bluck et al., 2020a; Hani et al., 2020; Pessa et al., 2021). We find a slope of $1.53$ at $\Sigma_{\rm gas}>10^{7.13\pm 0.02}$ for the rKS, with estimates in the literature ranging from $\sim 1.3-\sim 3.1$ To our knowledge, there are not comparable measurements of the total gas main sequence in the literature, but we note that the slopes we identify are all much smaller than estimates for the molecular gas main sequence (e.g., Pessa et al., 2021).
While quantitative comparisons of slopes are challenging, we can qualitatively compare our resutls for resolution dependence to other studies. We have shown that at $\Sigma_{*}<\Sigma_{*}^{\rm to}$, the rSFMS slopes we obtain with the ridge line method agree with each other within $2\sigma$ in almost all cases. The only exception is the res500 case, where the low-mass slope agrees with all others only at $3\sigma$ [^6] We note that a resolution of $500$ pc is close to the spatial resolution limit of TNG100, which has a gravitational softening length of $\sim 180$ pc. The resolution dependence of the rSFMS has also been explored by Hani et al. (2020) and Pessa et al. (2021). Hani et al. (2020) used ordinary least squares to identify the rSFMS for galaxies from the FIRE project and found that the slope became steeper as spatial resolution decreased. While we find that the rSFMS slope at low masses ($\gamma$) is roughly constant with changing spatial resolution, we find that the slope at high masses ($\gamma^{\prime}$) becomes steeper as spatial resolution decreases. The ‘spaxel’ samples used by Hani et al. (2020) appear to be biased toward stellar masses greater than the turnover masses we have identified ($\Sigma_{*}^{\rm to}$), which means the slopes they obtained would be biased toward the high-mass end of the relation. Thus, we find that our results are consistent with those of Hani et al. (2020). Pessa et al. (2021) fit the rSFMS, resolved molecular gas main sequence, and molecular rKS relation with a single power law and found that the slopes were independent of spatial resolution in a spaxel sample drawn from the Physics at High Angular resolutions in Nearby GalaxieS (PHANGS) survey (Leroy et al., 2021). For all three relations, the conditional means reported for the PHANGS data appear to deviate from an ideal power law.
IFS-capable telescopes are increasingly common, opening the door to more spatially resolved data for observed galaxies. Specific proposed and ongoing programs and instruments that will collect IFS data for galaxies include the Galaxy Assemby NIRSpec IFS survey (GA-NIFS; Perna, 2023) using JWST, the first generation Extremely Large Telescope instrument High Angular Resolution Monolithic Optical and Near-Infrared Integral field spectrograph (HARMONI; Thatte et al., 2020), the Atacama Large Millimetre Array Large Program [CII] Resolved ISM in Star-forming Galaxies (ALMA-CRISTAL; e.g., Li et al., 2024), and the Sloan Digital Sky Survey V Local Volume Mapper survey (LVM; Drory et al., 2024). The combination of these data sets and a measurement technique that is consistent across varying spatial resolutions would enable studies into the redshift evolution of the rSFMS and the other relations. Further, spatially resolved maps of SFR, stellar mass, and gas mass can and often have already been obtained from high-resolution simulations such as FIRE-2, Figuring Out Gas and Galaxies in Enzo (FOGGIE; e.g., Peeples et al., 2019), and the Assembling Galaxies of Resolved Anatomy project (AGORA; Kim et al., 2014). The multitude of possible sources for spatially resolved galaxy data, of which we have only named a selection, emphasizes the need for consistent, objective measurements of scaling relations.
The Kennicutt (1989) results were based on a sample of $15$ galaxies sampled with a spatial resolution of $1-5$ kpc. With a comparative wealth of data available now and in the coming years, it should be possible to constrain the dependence of $\Sigma_{\rm SFR}$ on $\Sigma_{*}$ and $\Sigma_{\rm gas}$ and explain deviations from these relations in terms of true stochasticity or other physical parameters such as metallicity, turbulence, and magnetic field strength. However, it is critical that the community come to a consensus on best practices for fitting scaling relationships and describing their scatter. As it is, estimates of the power law slopes in the literature result from a variety of different fitting methods which are often not described with sufficient detail for reproduction. This motivates a consistent reanalysis of existing data with clearly defined, physically-motivated methodology. Such a reanalysis could constrain the influence of resolution, SFR diagnostics and other factors on the resulting relationships. Our results indicate that the ridge line technique we present here is well suited for such a reanalysis but there are further improvements that can be made to improve the statistical rigor. We have published ScaleRPy as an open-source code so that further improvements can be made to the methodology as the community comes to a consensus on statistically rigorous best fitting practices (e.g., following Hogg et al., 2010) and physically-motivated functional forms for the star-forming scaling relationships. Simulations that reproduce spatially resolved galaxies, such as TNG100, are an excellent laboratory for testing fitting techniques because there are often known truth values in the form of the analytical models that describe star formation.
@@
In conclusion, we find that the ridge line technique offers significant benefits over standard least squares when it comes to fitting scaling relationships. For the rSFMS, fitting the ridge line with a double power law reveals that the slope at the low-mass end of the rSFMS is independent of spatial resolution and smoothing scale. This method enhances the comparability of rSFMS measurements because it does not require pre-selection of star-forming spaxels and parameterizes the location at which the rSFMS slope turns over. For all three spatially resolved scaling relationships we have explored here, the ridge line method results in significantly better fits than least squares. Thus, the ridge line technique is a promising method for making objective measurements of scaling relationships that can be compared across studies. We argue that such objective measurements are highly desirable if the physical processes that govern star formation are to be understood. The open-source ScaleRPy package facilitates the adoption and improvement of the method presented here for both global and resolved scaling relations.